# Optimizing Token Choice for Code Watermarking: A RL Approach

### v1

@@ Abstract
The need for detecting LLM-generated code necessitates watermarking systems capable of operating within its highly structured and syntactically constrained environment. To address this, we introduce CodeTracer, an innovative adaptive code watermarking framework underpinned by a novel reinforcement learning training paradigm. At its core, CodeTracer features a policy-driven approach that utilizes a parameterized model to intelligently bias token choices during next-token prediction. This strategy ensures that embedded watermarks maintain code functionality while exhibiting subtle yet statistically detectable deviations from typical token distributions. To facilitate policy learning, we devise a comprehensive reward system that seamlessly integrates execution feedback with watermark embedding signals, balancing process-level and outcome-level rewards. Additionally, we employ Gumbel Top-k reparameterization to enable gradient-based optimization of discrete watermarking decisions. Extensive comparative evaluations demonstrate CodeTracer’s significant superiority over state-of-the-art baselines in both watermark detectability and the preservation of generated code’s functionality.
@@ 1 Introduction
The unprecedented capabilities of large language models (LLMs) in code generation have introduced critical challenges for intellectual property protection and code attribution [27, 1]. As AI systems produce increasingly sophisticated code that is difficult to distinguish from human-written, the need for reliable code tracing approaches has become urgent [26]. Code watermarking, a technique historically vital for software protection, now serves as the cornerstone for identifying and verifying machine-generated code origins [9, 11, 4]. Unlike natural language, code presents unique watermarking challenges due to its strict structural and semantic requirements: even slight modifications can completely break functionality, a problem less prevalent in text watermarking [13].
LLM generation is an autoregressive process composed of sequential next-token predictions [20]. LLM watermarking biases these predictions toward token choices that subtly differ from normal selections, creating detectable statistical patterns. Many LLM watermarking schemes typically rely on a “green-red list” methodology [9]. This technique biases token generation toward specific green (preferred) tokens and suppressing red (avoided) tokens, creating detectable statistical patterns without significantly degrading output quality. For natural language, these vocabulary partitions can be determined randomly [9, 13] with minimal impact on coherence. Code watermarking faces unique challenges: random token selection often breaks functionality since certain tokens are syntactically essential, while uniform watermarking ignores structural constraints where some positions cannot tolerate modifications.
Therefore, we identify our core challenge: how to intelligently determine optimal watermark insertion points and select semantically reasonable token choices that maintaining statistical detectability patterns while preserve code functionality?
@@
Traditional code watermarking approaches suffer from significant limitations: they operate as post-generation transformations on completed code, making them incompatible with the causal generation process of LLMs. Additionally, these methods require labor-intensive, language-specific transformation rules that must be manually crafted for each programming language [8, 24, 14]. Recent efforts to incorporate watermarks during LLM code generation show promise but remain impractical in real-world scenarios, as they require access to supplementary prompts and model information to compute critical values like entropy during detection [13, 6]. We argue that an effective solution requires a model-based approach that learns the necessary programming knowledge during training, enabling intelligent watermarking decisions that adapt to syntactic and semantic constraints with minimal context.
In this work, we propose CodeTracer, an adaptive code watermarking framework built on a novel reinforcement learning (RL) paradigm. Our approach features a parameterized watermark model that intelligently identifies optimal positions and biases token selection during the code generation process. Reinforcement learning provides an ideal solution for our task as it eliminates the need for pre-existing watermarked code examples. This is crucial because creating such training examples presents a fundamental paradox: we cannot determine whether a code snippet is “watermarked” without first defining the watermark model itself.
Challenges with RL.
Designing a learnable code watermarking RL framework presents significant challenges. Two critical difficulties emerge in this process. First, effective reward mechanisms are fundamental to guiding the RL process, requiring careful integration of available feedback signals. Second, the watermarking procedure necessitates making discrete decisions about code tokens, creating a non-differentiable process that is inherently difficult to optimize. To address these challenges, we implement a dual-component reward system: execution feedback that penalizes functionally incorrect code, complemented by watermark embedding signals consisting of both immediate process rewards for successful watermarked token selection and statistical outcome rewards that employ metrics like z-score to comprehensively measure detectability. Additionally, we overcome the non-differentiability constraint through Gumbel Top-k reparameterization [22], enabling gradient-based optimization by allowing differentiable selection of tokens.
Contributions.
@@
Our work makes several key contributions: (i) We introduce CodeTracer, an adaptive watermarking framework that intelligently embeds watermarks during LLM code generation; (ii) We develop a novel RL pipeline for code watermarking, combines execution feedback with dual watermark signals, and enables differentiable optimization for the whole pipeline; and (iii) Our evaluations across multiple datasets demonstrate CodeTracer’s superior watermark detection capabilities while maintaining code functionality.
2 Related Works
LLM Watermarking.
@@ 2 Related Works
LLMs’ sophisticated content generation capabilities create challenges for authenticity verification [26, 19, 4]. Recent watermarking approaches embed signatures during the token sampling process by modifying logits or altering the sampling procedure [9, 11]. A notable example is the green-red watermarking scheme [9], which partitions vocabulary tokens into lists and biases selection toward “green” tokens, enabling statistical detection. [23] introduces a RL-based approach for watermark embedding but their co-training scheme may cause unexpected behavior for LLMs. However, these methods struggle with low-entropy conditions common in code generation [13].
Code Watermarking.
@@
Code watermarking imposes unique challenges due to programming languages’ strict syntactic and semantic requirements. Traditional code watermarking alters existing code through format changes or control flow restructuring [8, 17, 14, 24, 16, 4]. Recent LLM-focused approaches use entropy distributions [13, 15] or type predictors [6] to guide watermark insertion. However, these techniques typically require privileged access to LLM parameters, generation probabilities, or original prompts, limiting their practical deployment. CodeTracer instead adaptively embeds watermarks during generation without such prerequisites.
Rule-Based RL and GRPO.
Rule-Based Reinforcement Learning enhances LLMs by integrating explicit rules into the learning process, demonstrating reduced vulnerability to reward hacking [7, 12]. These approaches leverage simple, verifiable rules that ground model outputs, showing particular effectiveness on structured tasks like coding and mathematical reasoning [25]. Deepseek-R1 [7] exemplifies this approach through Group Relative Policy Optimization (GRPO), which significantly reduces computational overhead while improving complex reasoning capabilities. GRPO generates multiple candidate responses for a given prompt, then calculates advantages by normalizing rewards relative to group statistics, eliminating the need for trained critic models in value estimation. Our CodeTracer framework adopts this GRPO scheme with rule-based rewards for computational efficiency.
3 Preliminaries
LLM Watermarking.
In a code generation task, large language model receives a sequence of tokens $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ as the input prompt, and generates output $\mathbf{y}=[y_{1},y_{2},\ldots,y_{m}]$ including the code snippet. At every generation step $t$, the LLM $\pi_{\theta}$ receives the input prompt $\mathbf{x}$ and previous generated token to compute the logit vector $\mathbf{l}=[l_{1},l_{2},\ldots,l_{|\mathcal{V}|}]$ for token $t$, cross all tokens in the vocabulary $\mathcal{V}$, where each $l_{j}$ corresponds to token $v_{j}\in\mathcal{V}$.. The logit vector is then applied with a softmax opreator to transform it into a probability distribution where token at position $y_{t}$ is sampled from. To detect AI generated content, in-generation watermarking [9] adjust the original LLM generation policy by adding constant biases on a sublist of vocabulary that are selected by a pseudorandom function, resulting in a watermarked code sequence $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m^{\prime}}]$. The statistical difference in the token distribution within/out the sublists are then utlized to detect the AI-generated output.
Challenge of LLM Code Watermarking.
@@ 3 Preliminaries
However, LLM watermark faces serve performance degration in the code because of unique challenges distinct from natural text watermarking. Unlike natural language, code is highly structured with a delicate syntax where small modifications can drastically alter functionality or render the code inoperable. This inherent rigidity creates significant constraints for watermarking techniques. First, watermark position selection is particularly critical in code, as certain positions are immutable. Unlike natural text where minor alterations rarely affect meaning, code contains strict structural elements that cannot be modified. For example, altering the def keyword in a Python function definition (e.g., changing def process_data(data): to func process_data(data):) inherently breaks syntax. Any watermarking approach must identify and avoid these critical positions to maintain functional code. Second, watermark token choice must respect contextual constraints. Even at positions where modification is possible, the replacement tokens must be semantically and syntactically appropriate. For instance, if a model generates an assignment like status = "active" but a watermarking scheme forces a substitution of "active" with an incompatible token such as the keyword class, the resultant status = class would cause a syntax error. This contextual sensitivity significantly limits the vocabulary available for watermarking while preserving code functionality.
Why RL for $\pi_{\phi}(a|\mathbf{c})$ training?
Reinforcement learning provides an ideal framework for this task for three key reasons: (i) Learning complex rules implicitly: The diversity of programming languages makes manually defining watermarking rules impractical. RL can learn syntactic constraints across various code structures through interaction rather than exhaustive specification. (ii) Overcoming the absence of supervised data: SFT is impossible due to a circular dependency—we cannot create "gold standard" watermarked examples without first defining the watermark model itself. RL enables learning optimal strategies without requiring predefined examples. (iii) Leveraging clear feedback signals: Both watermark effectiveness and code functionality provide quantifiable rewards. Our system utilizes rule-based feedback mechanisms that have proven effective for code tasks [7, 12], making RL well-suited for optimizing within code generation constraints.
4 Learn to Watermark
We formulate the LLM code watermarking problem within a reinforcement learning (RL) framework. This section first introduces the watermark policy, then outlines the process of training this policy through RL, and finally addresses specific design challenges along with our proposed solutions.
4.1 Watermarked Policy
We focus on a single step in the token generation process, specifically one next token prediction. Watermark model $\pi_{\phi}(a|\mathbf{c})$ operates based on the current context $\mathbf{c}$ (comprising part of input $\mathbf{x}$ and previously generated tokens in $\mathbf{y}$ up to the current position with a fixed number of tokens). This model outputs an action $a=(w,G)$ consisting of two components: a binary decision $w\in\{0,1\}$ indicating whether to apply watermarking at this position, and a green token list $G\subset\mathcal{V}$ for preferred tokens.
The watermarked token generation process combines the original LLM $\pi_{\theta}$ with the watermark model $\pi_{\phi}$ and get watermarked policy $\pi_{\theta\oplus\phi}$, computing the modified logits as:
$\tilde{l}_{j}=l_{j}+w\cdot\delta\cdot\mathbbm{1}_{v_{j}\in G}$ (1)
where $\delta$ represents a bias added exclusively to green tokens when watermarking is active ($w=1$), and $\mathbbm{1}_{v_{j}\in G}$ is an indicator function equaling 1 if token $v_{j}\in G$ and 0 otherwise. The watermark strength is controlled by two constraints: the size of the green token set where $|G|=\gamma|\mathcal{V}|$ for a predefined parameter $\gamma\in(0,1)$, and the bias value $\delta$. The remaining tokens form a “red” group $R=\mathcal{V}\setminus G$. This policy-driven, selective approach dynamically controls both watermark placement and vocabulary partitioning used for biasing.
With this formulation, the policy strategically biases token choices through the green set $G$ and controls watermark application via the binary decision $w$. The core challenge lies in training the parameters $\phi$ of the watermark policy model. In the following subsection, we present the complete watermark pipeline for both watermark generation and detection using the combined policy $\pi_{\theta\oplus\phi}$, followed by our training methodology that optimizes the watermark policy $\pi_{\phi}$ in next section.
Generation
@@ 4.1 Watermarked Policy
During code generation, our framework CodeTracer applies the previously defined framework sequentially. The watermarked policy $\pi_{\theta\oplus\phi}(\mathbf{\tilde{y}}|\mathbf{x})$ takes exactly same input with LLM $\pi_{\theta}$ but generate the biased output sequences $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots]$ based on biased logits $\tilde{l}_{j}$ at each step.
Detection.
@@
For detection, CodeTracer analyzes a potentially corrupted sequence $\mathbf{s}=[s_{1},s_{2},\ldots,s_{T^{\prime}}]$ of length $T^{\prime}$. The watermark policy $\pi_{\phi}$ reconstructs its decisions independently, without requiring the LLM $\pi_{\theta}$. This reconstruction process determines two critical pieces of information for each token: first, whether watermarking was applied at that position ($w=1$ or $w=0$), and second, which tokens belonged to the green list $G$ if watermarking was applied. By applying this reconstruction, we identify the subset of tokens $s_{w=1}$ where watermarking was active, yielding a total count of watermarked positions $T=|\{s_{w=1}\}|$. Among these watermarked positions, we then count how many tokens fall into their respective green sets, denoted as $N_{G}=|\{s_{w=1,s\in G}\}|$. A one-proportion $z$-test assesses statistical significance:
@@
A significantly large positive $z$-score indicates watermark presence, demonstrating that watermarked tokens appear in their respective green sets more frequently than would occur by random chance.
Figure 1: CodeTracer: A framework for LLM code watermarking through selective token biasing. The diagram illustrates the end-to-end pipeline where a trainable watermark model works alongside an LLM to embed statistical patterns in code, with a reward system that optimizes for both code functionality and watermark detectability.

### v2

# Optimizing Token Choice for Code Watermarking: An RL Approach

@@ Abstract
Protecting intellectual property on LLM-generated code necessitates effective watermarking systems that can operate within code’s highly structured, syntactically constrained nature. In this work, we introduce CodeTracer, an innovative adaptive code watermarking framework underpinned by a novel reinforcement learning training paradigm. At its core, CodeTracer features a policy-driven approach that utilizes a parameterized model to intelligently bias token choices during next-token prediction. This strategy ensures that embedded watermarks maintain code functionality while exhibiting subtle yet statistically detectable deviations from typical token distributions. To facilitate policy learning, we devise a comprehensive reward system that seamlessly integrates execution feedback with watermark embedding signals, balancing process-level and outcome-level rewards. Additionally, we employ Gumbel Top-k reparameterization to enable gradient-based optimization of discrete watermarking decisions. Extensive comparative evaluations demonstrate CodeTracer’s significant superiority over state-of-the-art baselines in both watermark detectability and the preservation of generated code’s functionality. Our code is available at https://anonymous.4open.science/r/CodeTracer-B8EE.
@@ 1 Introduction
The unprecedented capabilities of large language models (LLMs) in code generation have introduced critical challenges for intellectual property protection and code attribution (Li et al., 2022; Achiam et al., 2023; Guo et al., 2024; Hui et al., 2024). As AI systems produce increasingly sophisticated code that is difficult to distinguish from human-written code, the need for reliable code tracing approaches has become urgent (Zhao et al., 2024; Wang et al., 2024). Traditional code watermarking approaches apply post-generation transformations to completed code, making them inherently incompatible with the autoregressive generation process of LLMs. Moreover, these methods require labor-intensive, language-specific transformation rules that must be manually crafted for each programming language (Hamilton & Danicic, 2011; Yang et al., 2023; Li et al., 2024).
Existing LLM watermarking approaches operate on the autoregressive generation process by biasing next-token predictions toward statistically detectable patterns (Radford et al., 2019; Kirchenbauer et al., 2023a). In natural language contexts, this approach succeeds because text generation is robust, as most positions admit multiple semantically valid token choices, allowing watermarking to be applied across all generated tokens (Kuditipudi et al., 2023; Dathathri et al., 2024; Liu & Bu, 2024).
@@
However, code watermarking presents distinct challenges that arise from the structural and semantic constraints inherent to programming languages. Code generation imposes two critical constraints. First, syntactic dependencies severely constrain the space of valid token choices, as certain tokens are syntactically mandatory and their modification results in compilation failures (Guan et al., 2024). Second, code positions exhibit heterogeneous sensitivity to modifications, where indiscriminate watermarking strategies fail to account for the varying tolerance to perturbations across different code locations (Lee et al., 2023). Recent efforts to incorporate watermarks during LLM code generation show promise but remain impractical in real-world scenarios, as they require access to supplementary prompts and model information to compute critical values like entropy during detection (Lee et al., 2023; Guan et al., 2024; Zhao et al., 2024). We argue that an effective solution requires a model-based approach that learns the necessary programming knowledge during training, enabling intelligent watermarking decisions that adapt to syntactic and semantic constraints with minimal context.
Consequently, we identify the core challenge in LLM code watermarking: how can we intelligently determine optimal watermark insertion points and select semantically reasonable token choices that maintain statistical detectability while preserving code functionality?
In this work, we propose CodeTracer, an adaptive code watermarking framework built upon a novel reinforcement learning (RL) training paradigm. CodeTracer employs a policy-driven approach that leverages a parameterized model to intelligently identify optimal insertion positions and guide token selection during the code generation process. The parameterized model collaborates with the original LLM to form a watermarked policy, where only the parameterized model undergoes optimization during training. At generation time, the LLM provides logits while the parameterized model intelligently determines watermark application and token selection decisions. To train the watermarked policy, we employ reinforcement learning through a carefully designed dual-component reward system. This system integrates two complementary feedback signals: execution feedback that penalizes functionally incorrect code, and watermark embedding signals that comprise both immediate process rewards for successful watermarked token selection and statistical outcome rewards employing metrics such as the z-score to comprehensively assess detectability performance. To enable end-to-end gradient-based optimization, we address the non-differentiability of discrete bias token selection by Gumbel Top-k reparameterization (Xie & Ermon, 2019) and Straight-Through Estimation (Bengio et al., 2013), achieving differentiable watermarked policy training.
@@
Contributions. Our work yields several key contributions: (i) We introduce CodeTracer, an adaptive watermarking framework that intelligently embeds watermarks during LLM code generation; (ii) We develop a novel RL pipeline for code watermarking training that combines execution feedback with dual watermark signals and enables differentiable optimization for the entire pipeline; and (iii) Empirically, we validate the effectiveness of CodeTracer through comprehensive evaluations, demonstrating its superior watermark detection capabilities while maintaining code functionality.
2 Related Work
@@
LLM Watermarking. Existing LLM watermarking approaches embed imperceptible signatures during token sampling by modifying logits or altering the sampling procedure (Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; Zhao et al., 2023; Christ et al., 2024; Dathathri et al., 2024). A prominent example is the green-red watermarking scheme (Kirchenbauer et al., 2023a; b), which partitions the vocabulary into “green” (preferred) and “red” (avoided) tokens, biasing generation toward green tokens while suppressing red ones, enabling statistical detection of watermarked content. Xu et al. (2024) propose a reinforcement learning approach for watermark embedding that requires training the LLM, which may cause unexpected behaviors for LLMs. Critically, these methods encounter difficulties in low-entropy scenarios typical of code generation (Lee et al., 2023).
@@
Code Watermarking. Code watermarking presents distinct challenges due to the strict syntactic and semantic constraints inherent in programming languages. Traditional approaches modify existing code through formatting changes or control flow restructuring (Hamilton & Danicic, 2011; Ma et al., 2019; Li et al., 2024; Yang et al., 2023; Liu et al., 2024; Dathathri et al., 2024). Recent LLM-focused methods leverage entropy distributions (Lee et al., 2023; Li et al., 2023) or type predictors (Guan et al., 2024) to guide watermark insertion. However, these techniques typically require privileged access to LLM parameters, generation probabilities, or original prompts during watermark detection, significantly limiting their practical deployment (Zhao et al., 2024). In contrast, CodeTracer adaptively embeds watermarks during generation without such prerequisites.
RLVR and GRPO. Reinforcement Learning with Verifiable Rewards (RLVR) enhances LLMs by integrating verifiable reward signals into the training process, demonstrating improved robustness against reward hacking and superior performance (Lambert et al., 2024; Guo et al., 2025; Team et al., 2025). However, RLVR is constrained by scarcity of reliably verifiable signals, with most applications limited to mathematical problems and code execution tasks. DeepSeek-R1 (Guo et al., 2025) combines verifiable rewards with Group Relative Policy Optimization (GRPO), which improves computational efficiency over traditional policy optimization methods. Code watermarking emerges as a natural fit for RLVR, as watermark detection provides unambiguous, token-level verification signals that can be efficiently computed. We exploit this natural alignment by employing GRPO with watermark-based verifiable rewards to achieve both computational efficiency and robust watermarking.
3 CodeTracer: A Policy-Driven Watermarking Framework
Problem Setting.
In code generation tasks, a large language model takes a sequence of input tokens representing the prompt $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ and generates an output sequence $\mathbf{y}=[y_{1},y_{2},\ldots,y_{m}]$ containing the generated code. At each generation step $t$, the LLM $\pi_{\theta}$ processes the input prompt $\mathbf{x}$ and previously generated tokens to compute a logit vector $\mathbf{l}=[l_{1},l_{2},\ldots,l_{|\mathcal{V}|}]$ over the entire vocabulary $\mathcal{V}$. Each value $l_{j}$ represents the model’s preference for token $v_{j}\in\mathcal{V}$. This logit vector is transformed into a probability distribution using the softmax function, from which token $y_{t}$ is sampled. To enable detection of LLM-generated code, in-generation watermarking techniques, such as the approach by Kirchenbauer et al. (2023a), modify the LLM’s original logits $\mathbf{l}$ to form watermarked logits $\mathbf{\tilde{l}}$. This modification is achieved by adding biases to the logits of a specific subset (green list) $G\subset\mathcal{V}$ of the vocabulary, selected using a pseudorandom function (PRF). This process yields a watermarked code sequence $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m^{\prime}}]$. The presence of watermark is then statistically inferred by analyzing frequency difference of tokens appearing within the biased vocabulary subset.
Challenges in LLM Code Watermarking.
@@ 3 CodeTracer: A Policy-Driven Watermarking Framework
LLM watermarking faces severe performance degradation in code generation due to unique challenges distinct from natural text watermarking. Unlike natural language, code is highly structured with precise syntax where small modifications can drastically alter functionality or render the code inoperable. This inherent rigidity creates significant constraints for watermarking techniques. First, watermark position selection is critical, as many positions in code are immutable. Unlike natural text where alterations rarely affect meaning, code contains structural elements that cannot be modified. For example, changing def to func in Python function definitions breaks syntax. Any watermarking approach must avoid these critical positions. Second, watermark token choice must respect contextual constraints. Even at modifiable positions, replacement tokens must maintain syntactic validity. For instance, if status = ‘‘active’’ is watermarked by replacing ‘‘active’’ with class, the result causes a syntax error. This contextual sensitivity severely limits the available vocabulary for watermarking. Consequently, effective code watermarking requires a deeper understanding of code structure and semantics.
3.1 CodeTracer
To address the aforementioned challenges, CodeTracer introduces a policy-driven framework that integrates a watermark model $\pi_{\phi}$ with an LLM $\pi_{\theta}$, yielding a composite watermarked policy $\pi_{\theta\oplus\phi}$ capable of generating watermarked code. At each generation step $t$, the watermark model $\pi_{\phi}(a|\mathbf{c})$ operates conditioned on the current context $\mathbf{c}$, defined as the concatenation of a segment of the input $\mathbf{x}$ and the sequence of previously generated tokens $\mathbf{y}_{<t}$ within a fixed-length window. The output of this policy is an action $a=(w,G)$, where $w\in\{0,1\}$ is a binary variable indicating whether to apply watermarking at the current position, and $G\subset\mathcal{V}$ represents a set of preferred “green” tokens.
The generation of a token is then achieved by sampling from a modified logit vector from the watermarked policy $\pi_{\theta\oplus\phi}$ as:
$\tilde{l}_{j}=l_{j}+w\cdot\delta\cdot\mathbbm{1}_{v_{j}\in G},$ (1)
where $\delta$ is a hyperparameter controlling the bias applied to the logits of tokens in the green list $G$ when watermarking is active ($w=1$), and $\mathbbm{1}_{v_{j}\in G}$ is an indicator function equaling 1 if token $v_{j}\in G$ and 0 otherwise. The watermark strength is governed by the size of the green token set, $|G|=\gamma|\mathcal{V}|$ for a predefined ratio $\gamma\in(0,1)$, and the bias magnitude $\delta$. The complement of the green set, $R=\mathcal{V}\setminus G$ constitutes the “red” token set. This policy-driven and context-aware approach allows for dynamic control over both the placement of the watermark and the vocabulary subset used for biasing the generation process. In essence, this formulation empowers the policy to strategically influence the token sampling process by preferentially selecting tokens from the green set $G$, while simultaneously regulating the watermark through the binary decision $w$.
@@ 3.1 CodeTracer
Consequently, during generation, CodeTracer applies the watermarked policy $\pi_{\theta\oplus\phi}(\mathbf{\tilde{y}}|\mathbf{x})$ that takes the same input as the LLM $\pi_{\theta}$ but generates biased outputs $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots]$ via modified logits $\tilde{l}_{j}$.
@@
Detection. During detection, given an output sequence $\mathbf{s}=[s_{1},s_{2},\ldots,s_{T^{\prime}}]$ of length $T^{\prime}$, CodeTracer reconstructs the watermarking decisions independently using only the watermark model $\pi_{\phi}$, without requiring access to the LLM $\pi_{\theta}$. This reconstruction determines, for each token, whether watermarking was applied ($w=1$ or $w=0$) and, if so, the composition of the corresponding green token set $G$. We identify the subset of tokens where watermarking was active, denoted as $\{s_{w=1}\}$, and count the total number of watermarked positions $T=|\{s_{w=1}\}|$. We then count how many of watermarked tokens belong to their respective reconstructed green sets, denoted as $N_{G}=|\{s\mathrel{\mathop{\ordinarycolon}}s\in\{s_{w=1}\}\land s\in G\}|$. To assess the statistical significance, we employ a one-proportion $z$-test for the proportion of watermarked tokens belonging to their predicted green sets:
@@
Under the null hypothesis of no watermarking, we expect $N_{G}$ to follow a binomial distribution with success probability $\gamma$. A sufficiently large positive $z$-score provides strong statistical evidence for watermark presence, indicating that watermarked tokens appear in their predicted green sets with frequency significantly higher than the expected random baseline of $\gamma$. The complete algorithms for watermark generation and detection are formalized in Algorithm 1 and Algorithm 2, respectively.
Figure 1: CodeTracer: A framework for LLM code watermarking through selective token biasing. The diagram shows our end-to-end pipeline where a trainable watermark model collaborates with an LLM to embed detectable statistical patterns in generated code. A reward system optimizes the dual objectives of preserving code functionality while maximizing watermark detectability. The watermark model operates as a plug-in module, enabling deployment beyond those used during training. Importantly, watermark detection requires only the watermark model, not the original LLM.
4 Learning to Watermark in CodeTracer
We formulate training of watermarked policy $\pi_{\theta\oplus\phi}$ within a reinforcement learning framework. This section first gives an overview of the training pipeline, then discusses the reason to use RL for policy learning, and finally addresses specific design challenges along with our proposed approach.